A simple random sample of 400 persons is taken to estimate the percentage of women in the Boston Teacher's Union whose total membership is 8000. Of those surveyed, 268 teachers are females. We can be 95 percent confident that between __ percent and ___ percent of the teachers in the union are females.
Answer by Boreal(15235)

268/400=72%
95% interval has width 1.96*sqrt(.5*.5/400) or 0.049
the exact interval is (0.671, 0.779)
estimate by margin of error is 1/sqrt(n)=1/20=0.05, and get (0.67, 0.77)